In a certain clothing store, 3 shirts and 5 ties cost $60, and 2 shirts and 3 ties cost $39. What is the cost of each shirt?
If s = the cost of a shirt and t = the cost of a tie, which system of equations represents the problem?
A.)3t + 5s = 60 and 2t + 3s = 39
B.)3s + 5t = 60 and 2s + 3t = 39
C.)s + t = 60 and 5s + 8t = 99

Answer by ikleyn(52788)

.
In a certain clothing store, 3 shirts and 5 ties cost $60, and 2 shirts and 3 ties cost $39. What is the cost of each shirt?
If s = the cost of a shirt and t = the cost of a tie, which system of equations represents the problem?
A.) 3t + 5s = 60 and 2t + 3s = 39
B.) 3s + 5t = 60 and 2s + 3t = 39
C.) s + t = 60 and 5s + 8t = 99
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Can you read and differentiate letters ans words? And the numbers, too ?